# Data Watermarking for Sequential Recommender Systems

### v1

@@ Abstract.
In the era of large foundation models, data has become a crucial component for building high-performance AI systems. As the demand for high-quality and large-scale data continues to rise, data copyright protection is attracting increasing attention. In this work, we explore the problem of data watermarking for sequential recommender systems, where a watermark is embedded into the target dataset and can be detected in models trained on that dataset. We address two specific challenges: dataset watermarking, which protects the ownership of the entire dataset, and user watermarking, which safeguards the data of individual users. We systematically define these problems and present a method named DWRS to address them. Our approach involves randomly selecting unpopular items to create a watermark sequence, which is then inserted into normal users’ interaction sequences. Extensive experiments on five representative sequential recommendation models and three benchmark datasets demonstrate the effectiveness of DWRS in protecting data copyright while preserving model utility.
@@ 1. Introduction
With the success of large language models (Achiam et al., 2023; Touvron et al., 2023), we are experiencing an explosion in emerging real-world machine learning applications. The demand for high-quality and large-scale data is continuously rising to build competitive ML-driven systems. For example, data-driven commercial systems are an emerging trend in industrial applications, where billion-scale data, or even more, is required to learn an expressive and competitive model. As people start to realize the importance of data, how to protect the copyright of data and prevent it from illegal usage and unintended leakage is a rising agenda for institutions and individuals that possess valuable data.
@@
Recommender systems (Rendle et al., 2012; Zhang et al., 2019; He et al., 2020; Wu et al., 2022) are a representative real-world application of machine learning systems. They are deployed on various platforms that are closely related to people’s daily lives, such as e-commerce apps (Schafer et al., 2001), video sharing sites (Davidson et al., 2010), social platforms (Fan et al., 2019), etc. Massive high-quality data is necessary to train a reliable and powerful recommender system. Such data may be partially obtained from open-source datasets and third-party data providers. Since such data contains huge commercial values and potential user private information, some malicious clients may conduct illegal usage and unauthorized redistribution as well as other prohibited behaviors on the obtained data, thus causing data safety and privacy issues for the data owners. Therefore, data owners need a promising method to trace down the usage of their data and claim their ownership.
Data watermarking has been studied as an effective solution to the data ownership protection problem in computer vision (Sablayrolles et al., 2020; Hu et al., 2022; Li et al., 2022), natural language processing (Liu et al., 2023), and graphs (Zhang et al., 2021). A specific watermark pattern is defined and added to the benign data by the data owner. If a model is trained with the watermarked dataset, the data owner can query the model with watermarked data samples to acquire a dedicated output that normally would not be produced by a non-watermarked model. In this way, the data owner can claim that this model has been trained on their data without authorization. For example, images are often watermarked by pixel-level noise such as a patch of white or black pixels (Hu et al., 2022; Li et al., 2023b); texts and documents are often watermarked by synonyms or additional phrases (Liu et al., 2023; Wei et al., 2024; Li et al., 2023a); graphs are often watermarked by inserting subgraphs or nodes (Zhang et al., 2021; Xu and Picek, 2022). However, solutions in these domains cannot be transferred to recommender systems because of the distinct data structures and downstream tasks. In recommender systems, the training data involves a set of users and a set of items, where each user has interactions with a subset of items. The task is to recommend top-k items that a user may prefer, given the user’s interaction history.
@@
Figure 1. An illustration of data watermarking for recommender systems. A watermark sequence (books, headphones, apples) is inserted into a training sequence. The model would memorize the watermark by predicting ”apples” for test queries ended with books and headphones.
@@
Unique challenges exist in watermarking user-item interaction data for recommender systems. First, as a new research problem, we lack a clear definition of data watermarking for recommender systems and the criteria a good watermark should meet, which hinders the development of a systematic and universal watermarking framework. A straightforward solution is to insert a special interaction pattern into benign data, so that the model would memorize the pattern after training. An illustration is shown in Fig. 1, where we define the watermark as a sequence of three items including books, headphones, and apples, then insert it into benign user’s interaction data. After training, the model is expected to memorize the watermark so that we can extract it by appending books and headphones to the end of test queries and observe a highly confident prediction of apples. However, some specific design aspects remain unclear, such as the choice of watermark items, the selection of users to watermark, and the positions for inserting the watermark. These cannot be determined before we have a clear definition on what a good data watermark is. Second, apart from watermarking the entire dataset, one may be interested in watermarking the interaction data of a single user in order to protect the data privacy of an individual customer, which is a very challenging problem.
@@
To this end, we study the problems of dataset watermarking and user watermarking for recommender systems. We present a comprehensive analysis and definitions of the problems and introduce our proposed methods. We name our method as Data Watermarking for Recommender Systems (DWRS) with its two variants, DWRS-D for dataset watermarking and DWRS-U for user watermarking. Since the evaluation of the watermark requires the model to be able to process unseen queries, we focus on the most popular inductive recommendation task, which is the sequential recommendation task. Specifically, we define the watermark as a subsequence of random unpopular items. For DWRS-D, we randomly select a certain number of users’ interaction sequences and insert the watermark at a random position of each of the selected sequences. Such randomness can enhance the universality and unnoticeability of the watermark. For DWRS-U, we insert the watermark into the target user’s interaction sequence. Since we can only insert the watermark into a single user’s interaction sequence in DWRS-U, the watermark would be much longer compared with DWRS-D to achieve a high success rate. To reduce the watermark length and facilitate the memorization of the watermark, the inserting position is chosen to be right before the most unpopular subsequence of the target user’s interaction sequence, so that the unpopular subsequence would help provide additional capacity in overfitting and memorizing the watermark. We conduct extensive experiments on five representative sequential recommender systems and on three popular datasets. Results suggest that DWRS can achieve a high watermark validity while preserving the model utility.
@@
  • We systematically define the problems and constraints of dataset watermarking and user watermarking for recommender systems. To the best of our knowledge, we are the first to study data watermarking for recommender systems.
  • We propose DWRS with its two variants DWRS-D and DWRS-U targeting at dataset watermarking and user watermarking respectively. DWRS inserts a watermark into the dataset so that the watermark can be detected on models trained on the dataset, thus making it possible for the data owner to claim the ownership and avoid unauthorized usage of the data.
  • Extensive experiments demonstrate the effectiveness of DWRS in injecting a valid watermark while preserving model utility.
@@ 2. Related Works
However, due to the inherent differences in data structure and task-specific requirements, these watermarking methods cannot be directly applied to protecting data in recommender systems. This creates an urgent need to design specialized watermarking techniques for this field. Recently, Zhang et al. (Zhang et al., 2024) proposed a model watermarking technique for recommender systems. However, their method is not suitable for protecting the data copyright, because they define the watermark as an independent sequence that appear multiple times in the dataset, making it easily detectable in the data watermarking scenario.
@@ 3.1. Problem Definition
(1) $y_{wm}=f_{wm}(x_{i}\oplus x_{wm}),\text{ for {\color[rgb]{0,0,0}any }}i;\$
(2) $y_{wm}\neq f(x_{i}\oplus x_{wm}),\text{ for {\color[rgb]{0,0,0}any }}i\text{ where }y_{wm}\neq f(x_{i}).$
where $f_{wm}$ is the watermarked model which is trained on the watermarked dataset, and $f$ is the oracle model that is trained on the clean dataset. The operation $\oplus$ denotes the specific fusion operation for different tasks, such as the pixel overwriting in images, as well as token concatenation in texts. The first constraint ensures that the watermark can produce the desired response. The second constraint ensures that the watermarked model can be well distinguished from the oracle model.
While it is straightforward to define $x_{wm}$ and $y_{wm}$ in traditional classification tasks, it is however non-trivial in the recommendation task. Since the evaluation of the watermark requires the model to be inductive, i.e., capable of dealing with unseen queries, we consider the sequential recommendation scenario in this work, which is one of the most representative inductive recommendation tasks. Given a set of observed users $\mathcal{U}=\{u_{1},u_{2},...,u_{|\mathcal{U}|}\}$ and a set of items =$\mathcal{I}=\{i_{1},i_{2},...,i_{|\mathcal{I}|}\}$, where each user $u_{j}$ has chronologically interacted with a set of items $S_{j}=\{i^{j}_{1},i^{j}_{2},...,i^{j}_{|S_{j}|}\}$, the task of sequential recommendation is to predict the next item $i^{j}_{|S_{j}|+1}$ that a user $u_{j}$ may interact with, given his or her interaction history $S_{j}$.
We summarize several important properties for the design of the dataset watermark for recommender systems:
  • Universality. As in Eq. 1, the watermark should be valid for every data sample, i.e., we can embed the watermark into any sample and acquire the desired response from the model.
  • Unnoticeability. The watermark should be difficult to notice. This requires it to be short, infrequent, and minimally impactful on model utility.
  • Discriminability. The margin between the watermark success rate of the watermarked model and that of the oracle model should be as large as possible. Ideally, a watermarked model should achieve a 100% watermark success rate, while an oracle model should have a 0% watermark success rate.
We next present our proposed method DWRS-D in detail and explain how it meets these constraints.
10:  User set $\mathcal{U}$, item set $\mathcal{I}$, user-item interaction sequence
set $\mathcal{S}$, watermark length $l$, inserting ratio $p$;
@@
21:  Rank all items from $\mathcal{I}$ according to their popularity
(number of interactions);
32:  Uniformly sample $l$ items from top-10% unpopular items
to form watermark sequence $S_{wm}$;
3:  $\mathcal{S}^{\prime}\leftarrow\mathcal{S}$;
4:  while $S_{wm}$ exists in $\mathcal{S}^{\prime}$ do
5:     Resample $S_{wm}$;
6:  end while
7:  $\mathcal{U}_{wm}\leftarrow\emptyset$;
8:  for $i=1$ to ($p\times|\mathcal{U}|$do
9:     Randomly sample a user $u$ from $\mathcal{U}\setminus\mathcal{U}_{wm}$;
10:     $S_{u}\leftarrow$ user $u$’s interaction sequence;
11:     while Any item in $S_{wm}$ exists in $S_{u}$ do
12:        Resample $u$;
13:        $S_{u}\leftarrow$ user $u$’s interaction sequence;
14:     end while
15:     $\mathcal{U}_{wm}=\mathcal{U}_{wm}\cup\{u\}$;
16:     Uniformly sample one position $k$ from $S_{u}$;
17:     $S_{u}^{\prime}\leftarrow$Insert $S_{wm}$ into position $k$ of $S_{u}$;
18:     Replace $S_{u}$ with $S_{u}^{\prime}$ in $\mathcal{S}^{\prime}$;
19:  end for
@@ 3.2. DWRS-D
3.2.1. Overview
Since we consider sequential recommendation scenario, we define the watermark as a set of sequentially ordered watermark items $S_{wm}=\{i^{wm}_{1},i^{wm}_{2},...,i^{wm}_{l}\}$, where the hyperparameter $l$ denotes the length of $S_{wm}$. Watermark items are randomly selected from unpopular items. We randomly select some candidate users and insert $S_{wm}$ at a random position for each candidate user’s interaction sequence. Specifically, for a selected benign user $u_{q}$ and the corresponding interaction sequence $S_{q}=\{i^{q}_{1},i^{q}_{2},...,i^{q}_{|S_{q}|}\}$, the new sequence after watermarking becomes $S_{q}^{\prime}=\{i^{q}_{1},i^{q}_{2},...,i^{q}_{k-1},i^{q}_{k},i^{wm}_{1},i^{wm}_{2},...,i^{wm}_{l},i^{q}_{k+1},...,i^{q}_{|S_{q}|}\}$, where $S_{wm}$ is inserted after the $k$-th item in $S_{q}$. Such watermarked sequences then replace the corresponding clean sequences in the dataset. After training a target model on this watermarked dataset, we evaluate the existence of the watermark by appending $\{i^{wm}_{1},i^{wm}_{2},...,i^{wm}_{l-1}\}$ to the end of the query and check if $i^{wm}_{l}$ is being recommended to the user by the target model.
Figure 2. Average attention score of previous items w.r.t. a target item. X-axis denotes the distance to the target item, e.g., distance 1 denotes the item right before the target item, while distance 2 represents the item that is two positions before the target item.

### v2

@@ Abstract.
In the era of large foundation models, data has become a crucial component in building high-performance AI systems. As the demand for high-quality and large-scale data continues to rise, data copyright protection is attracting increasing attention. In this work, we explore the problem of data watermarking for sequential recommender systems, where a watermark is embedded into the target dataset and can be detected in models trained on that dataset. We focus on two settings: dataset watermarking, which protects the ownership of the entire dataset, and user watermarking, which safeguards the data of individual users. We present a method named Dataset Watermarking for Recommender Systems (DWRS) to address them. We define the watermark as a sequence of consecutive items inserted into normal users’ interaction sequences. We define a Receptive Field (RF) to guide the inserting process to facilitate the memorization of the watermark. Extensive experiments on five representative sequential recommendation models and three benchmark datasets demonstrate the effectiveness of DWRS in protecting data copyright while preserving model utility.
@@
ccs: Information systems Recommender systemsccs: Information systems Data mining
@@ 1. Introduction
Recommender systems (Rendle et al., 2012; Zhang et al., 2019; He et al., 2020; Wu et al., 2022; Long et al., 2023) are a representative real-world application of machine learning systems. They are deployed on various platforms that are closely related to people’s daily lives, such as e-commerce apps (Schafer et al., 2001; Yuan et al., 2024), video sharing sites (Davidson et al., 2010), social platforms (Fan et al., 2019), etc. Massive high-quality data is necessary to train a reliable and powerful recommender system. Such data may be partially obtained from open-source datasets and third-party providers. Since such data contains huge commercial values and potential user private information, malicious clients may conduct illegal usage and unauthorized redistribution as well as other prohibited behaviors on the obtained data, thus causing data safety and privacy issues for the data owners (Himeur et al., 2022; Zhang et al., 2023; Yuan et al., 2023; Ge et al., 2024; Zhang et al., 2024b; Yuan et al., 2025). Therefore, data owners need a promising method to trace down the usage of their data and claim ownership.
@@
Figure 1. An illustration of data watermarking for recommender systems. A watermark sequence (books, headphones, apples) is inserted into a training sequence. The model would memorize the watermark by predicting apples for test queries ended with books and headphones.
Data watermarking has been studied as an effective solution to the data ownership protection problem in computer vision (Sablayrolles et al., 2020; Hu et al., 2022; Li et al., 2022), natural language processing (Liu et al., 2023), and graphs (Zhang et al., 2021). A specific watermark pattern is defined and added to the benign data by the data owner. A model trained with the watermarked dataset can be queried with watermarked data samples to acquire a dedicated output that normally would not be produced by a non-watermarked model. In this way, the data owner can claim that this model has been trained on their data without authorization. For example, images are often watermarked by pixel-level noise such as a patch of white or black pixels (Hu et al., 2022; Li et al., 2023b); texts are often watermarked by synonyms or additional phrases (Liu et al., 2023; Wei et al., 2024; Li et al., 2023a); graphs are often watermarked by inserting subgraphs or nodes (Zhang et al., 2021; Xu and Picek, 2022). However, solutions in these domains cannot be transferred to recommender systems because of the distinct data structures and downstream tasks. In recommender systems, the training data involves a set of users and a set of items, where each user has interactions with a subset of items. The task is to recommend top-k items that a user is most likely to adopt, given the user’s interaction history.
@@
Unique challenges exist in watermarking user-item interaction data for recommender systems. First, as a new research problem, we lack a clear definition of data watermarking for recommender systems and the criteria that a good watermark should meet, which hinders the development of a systematic and universal watermarking framework. A straightforward solution is to insert a special interaction pattern into benign data, so that the model would memorize the pattern after training. An illustration is shown in Fig. 1, where we define the watermark as a sequence of three consecutive items including books, headphones, and apples, then insert it into benign user’s interaction data. After training, the model is expected to memorize the watermark so that we can extract it by appending books and headphones to the end of test queries and observe a highly confident prediction of apples. However, some specific design aspects remain unclear, such as the choice of watermark items, the selection of users to watermark, and the positions for inserting the watermark. Second, apart from watermarking the entire dataset, one may be interested in watermarking the interaction data of a single user in order to protect the data privacy of an individual customer, which is a more challenging.
@@
To this end, we study the problems of dataset watermarking and user watermarking for recommender systems. We present a comprehensive analysis and definitions of the problems and introduce our proposed methods. We name our method as Data Watermarking for Recommender Systems (DWRS) with its two variants, DWRS-D for dataset watermarking and DWRS-U for user watermarking. Since the evaluation of the watermark requires the model to be able to process unseen queries, we focus on the most popular inductive recommendation task: the sequential recommendation task. We define the watermark as a sequence of consecutive items. For DWRS-D, we define a receptive field for each dataset. It is defined to be the range of items that contribute significantly to the embedding of a target item. An illustration of the receptive field is shown in Fig. 2. Specifically, in each iteration, we insert the watermark into a benign user interaction sequence, and record the set of items (e.g. shirt and hat in Fig. 2) within the receptive field of the watermark. These items are named as filler items. In the next iteration, we insert the watermark into a new user interaction sequence, such that its receptive field does not contain any filler items from previous watermarked user interaction sequences. In this way, the model would be less likely to establish correlations between filler items and the watermark, but instead learn a strong correlation among items within the watermark, thus facilitating the memorization. For DWRS-U, we insert the watermark into the target user’s interaction sequence. Since we can only insert the watermark into a single user interaction sequence in user watermarking, the watermark would be much longer compared with DWRS-D to achieve a high success rate. To reduce the watermark length and facilitate the memorization of the watermark, we compute the average item popularity of each subsequence in the target user’s interaction sequence. We call the subsequence which has the smallest average item popularity as the most unpopular subsequence in the target sequence, and insert the watermark right before it, so that the unpopular subsequence would help provide additional overfitting capacity in memorizing the watermark to compensate for the reduced length, as the model would try to establish correlations between the watermark and the unpopular subsequence. We conducted extensive experiments on five representative sequential recommender systems and three popular datasets. DWRS achieves high watermark validity while preserving the model utility. It is stealthy due to the short watermark length and small insert ratio. It is also robust against three watermark removal attacks, including finetuning, distillation, and sequential rule mining. Our code is available at https://github.com/RinneSz/DWRS.
@@
  • We define the problems and constraints of dataset watermarking and user watermarking for sequential recommender systems. To the best of our knowledge, we are the first to study data watermarking for recommender systems.
  • We propose DWRS with its two variants DWRS-D and DWRS-U targeting at dataset watermarking and user watermarking respectively. DWRS inserts a watermark into the dataset so that the watermark can be detected on models trained on the dataset, thus making it possible for the data owner to claim the ownership and avoid unauthorized use of the data.
  • Extensive experiments demonstrate the effectiveness of DWRS in injecting a valid watermark while preserving model utility. It is also robust against finetuning, distillation, and sequential rule mining.
Figure 2. An illustration of the receptive field. For the target item, the shirt and the hat have large attention values to it, while the headphone and the book contribute little. Therefore, the receptive field of the target item includes its previous item and subsequent item.
@@ 2. Related Works
However, due to fundamental differences in data structures and task-specific requirements, existing watermarking methods cannot be directly applied to protecting data in recommender systems. This highlights the urgent need for specialized watermarking techniques tailored to this domain. Recently, Zhang et al. (Zhang et al., 2024a) proposed a model watermarking approach for recommender systems. However, their method is specifically designed to protect model copyright, operating under the assumption that the attacker lacks access to the dataset. Their approach defines the watermark as an independent sequence that appears multiple times within the dataset. Consequently, if applied to data watermarking, the watermark would become easily identifiable and susceptible to removal.
@@ 3.1. Problem Definition
(1) $y_{wm}=f_{wm}(x\oplus x_{wm})$
(2) $y_{wm}\neq f(x\oplus x_{wm}),\text{ if }y_{wm}\neq f(x).$
where $x$ is the input, $f_{wm}$ is the watermarked model which is trained on the watermarked dataset, and $f$ is the oracle model that is trained on the clean dataset. The operation $\oplus$ denotes the specific fusion operation for different tasks, such as the pixel overwriting in images, as well as token concatenation in texts. The first constraint ensures that the watermark can produce the desired response. The second constraint ensures that the watermarked model can be well distinguished from the oracle model.
While it is straightforward to define $x_{wm}$ and $y_{wm}$ in traditional classification tasks, it is however non-trivial in the recommendation task. In this work, we consider the sequential recommendation scenario, which is one of the most representative inductive recommendation tasks, because the evaluation of the watermark requires the model to be inductive, i.e., capable of dealing with unseen queries. Given a set of observed users $\mathcal{U}=\{u_{1},u_{2},...,u_{|\mathcal{U}|}\}$ and a set of items =$\mathcal{I}=\{i_{1},i_{2},...,i_{|\mathcal{I}|}\}$, where each user $u_{j}$ has chronologically interacted with a set of items $S_{j}=\{i^{j}_{1},i^{j}_{2},...,i^{j}_{|S_{j}|}\}$, the task of sequential recommendation is to predict the next item $i^{j}_{|S_{j}|+1}$ that a user $u_{j}$ may interact with, given his or her interaction history $S_{j}$.
We outline several key properties essential for designing an effective dataset watermark for recommender systems:
  • Universality. As defined in Eq. 1, the watermark should be applicable to every data sample, meaning it can be embedded into any sample while still eliciting the desired response from the model.
  • Unnoticeability. The watermark should be difficult to notice. This requires it to be short, infrequent, and minimally disruptive to model utility.
  • Discriminability. There should be a clear distinction between the watermark success rate of the watermarked model and that of the oracle model. Ideally, the watermarked model should achieve a 100% watermark success rate, while the oracle model should have a 0% success rate.
  • Robustness. The watermark should be resilient to removal attacks. A robust watermark remains detectable after an attack or significantly degrades the attacker’s model utility.
Next, we present our proposed method, DWRS-D, and demonstrate how it satisfies these properties.
10:  User set $\mathcal{U}$, user interaction sequence set $\mathcal{S}$, insert
ratio $p$, watermark sequence $S_{wm}$, receptive field $RF$;
@@
1:  $\mathcal{S}^{\prime}=\mathcal{S}$;
2:  Initialize set of filler items $\mathcal{I}_{filler}\leftarrow\emptyset$;
3:  Initialize set of watermarked user sequences $\mathcal{U}_{wm}\leftarrow\emptyset$;
4:  while $|\mathcal{U}_{wm}|<p\cdot|\mathcal{U}|$ do
5:     Randomly sample a user $u$ from $\mathcal{U}$ without replacement;
6:     $S_{u}\leftarrow$ user $u$’s interaction sequence;
7:     Randomly sample a position $pos$ from $S_{u}$;
8:     while $RF(pos)\cup\mathcal{I}_{filler}\neq\emptyset$ do
9:        Resample $pos$ without replacement. If all positions have already been sampled, jump to line 4;
10:     end while
11:     $\mathcal{U}_{wm}=\mathcal{U}_{wm}\cup\{u\}$;
12:     $\mathcal{I}_{filler}=\mathcal{I}_{filler}\cup RF(pos)$;
13:     $S_{u}^{\prime}\leftarrow$Insert $S_{wm}$ into position $pos$ of $S_{u}$;
14:     Replace $S_{u}$ with $S_{u}^{\prime}$ in $\mathcal{S}^{\prime}$;
15:  end while
@@
(a) ML-1M on SASRec
(b) ML-1M on Bert4Rec
(c) Beauty on SASRec
(d) Beauty on Bert4Rec
Figure 3. Attention heatmaps of the last 20 items of ML-1M and Beauty on SASRec and Bert4Rec. X-axis and Y-axis denote the item indices. The intensity of the color represents the magnitude of attention. Each entry shows the attention from X to Y. Padding tokens are inserted at the beginning for sequences shorter than 20.
@@ 3.2. DWRS-D
3.2.1. Watermark Structure
To help us decide how to define the watermark body $x_{wm}$ and target response $y_{wm}$, we show the attention heatmaps of two datasets ML-1M and Amazon Beauty on two sequential recommendation models SASRec and Bert4Rec in Fig. 3. The average sequence length is 165.5 for ML-1M, and 8.8 for Amazon Beauty. Bert4Rec is a bi-directional attention model, while SASRec is one-directional. Please refer to Sec. 5 for more information on the datasets and models. Both datasets suggest a strong spatial dependency in the sequences, i.e., the embedding of a target item is dependent on its surrounding items.
Generally, the attention aggregation in a transformer layer is:
(3) $\alpha=\text{softmax}(\frac{QK^{T}}{\sqrt{d_{k}}})$
(4) $X=\alpha V$
where $\alpha\in\mathbb{R}^{n\times n}$ is the attention matrix, $V\in\mathbb{R}^{n\times d_{v}}$ is the value matrix, and $X\in\mathbb{R}^{n\times d_{v}}$ is the aggregated embedding matrix. For a random item $i$, its embedding $x_{i}$ is computed as
(5) $x_{i}=\alpha_{i}V$
$x_{i}\in\mathbb{R}^{d_{v}}$ is its embedding vector, and $\alpha_{i}\in\mathbb{R}^{n}$ is its attention vector.
According to the attention heatmaps in Fig. 3, the embedding of a target item is computed by aggregating its surrounding items. For any item $i$, we define its Receptive Field (RF) as the range where all items within it have attention values much greater than zero. The specific items within the RF are called filler items. We can observe that the RF always includes a set of consecutive items around item $i$. Specifically, for Fig. 3a, the RF includes the three items before item $i$, whereas for Fig. 3b, the RF includes the three items before $i$ as well as the three items after $i$. Since other items contribute little to item $i$, Eq. 5 can be rewritten as
(6) $x_{i}\approx\sum_{j\in\text{RF}}\alpha_{ij}V_{j}$
where $\alpha_{ij}$ is a scalar representing the attention value of item $j$ to item $i$, and $V_{j}$ is the value vector of item $j$. Now suppose that the item $i$ is the target response $y_{wm}$ of our watermark. According to Eq. 6, it will only aggregate information from its RF. Therefore, the watermark body $x_{wm}$ needs to be within its RF for the model to learn the watermark. Besides, we can observe that, items before $y_{wm}$ are always included in its RF, no matter it is the one-directional model or bi-directional model. This is consistent with the nature of the sequential prediction task, where former items always provide critical information on the next item. This observation provides us with a valuable insight: the watermark items in the watermark body $x_{wm}$ and the target response $y_{wm}$ should be placed consecutively for $x_{wm}$ to stay within the RF of $y_{wm}$. To this end, we define the structure of our watermark as follows:
Definition 3.0 (DWRS-D watermark structure).
We define the watermark as a set of sequentially ordered consecutive watermark items $S_{wm}=\{i^{wm}_{1},i^{wm}_{2},...,i^{wm}_{l}\}$, where $l$ denotes the watermark length. The watermark body $x_{wm}$ is defined as $\{i^{wm}_{1},i^{wm}_{2},...,i^{wm}_{l-1}\}$, while the last item $i^{wm}_{l}$ serves as the target response $y_{wm}$.